Nestle Frozen Yogurt shops have enjoyed rapid growth in Karachi in recent years. From the analysis of Nestle various outlets, it was found that the demand curve follows this pattern.

         Q= 200 – 100P + 120i + 32T -250 Ac + 400 AN

Where

Q= Number of cups served per week

 P= Average price paid for each cup

 i= Per capita income in the given market (thousands)

 T= Average outdoor temperature

 Ac= Competition’s monthly advertising expenditures (thousands)

AN = Nestle own monthly advertising expenditure (thousands)


1
Expert's answer
2022-03-28T14:06:59-0400

One of the outlets has the following conditions;

P = 1.50, I= 10, T= 30 Ac = 15, An =10

a) Estimate the number of cups served per week by this outlet also determine the outlets demand curve.

"Q= 200-100P+120i+32T-250Ac+400An"

"Q= 200-100(1.5)+120(10)+32(30)-250(15)+400(10)"

"Q= 2460" cups per week.


Also determine the outlets demand curve

"Q= 200-100P+120i+32T-250Ac+400An"

"Q= 200-100(P)+120(10)+32(30)-250(15)+400(10)"

"Q= 200-100(P)+1200+960-3750+4000"

"Q= 2610-100P"